# On Sybil-Proof Mechanisms

### v4

@@ Appendix E Proof of Theorem 3
If $v$ is strictly increasing in $\mathbb{R}_{>0}\times(0,1]$ and twice differential function with $v(\theta,0)=0$ and $v(0,x)=0$ and we have type-convexity and superadditive interim utility then:
@@
(2) Fix $x^{\prime}>x$ and define $\phi(\theta):=v(\theta,x^{\prime})-v(\theta,x)$. Then $\phi^{\prime}(\theta)=\partial_{1}v(\theta,x^{\prime})-\partial_{1}v(\theta,x^{\prime})$. Because $v(\theta,0)=0$ and $x\mapsto v(\theta,x)$ is superadditive, it is non-decreasing on $(0,1]$; hence $\phi^{\prime}(\theta)\geq 0$ for all $\theta>0$. Thus $\phi$ is non-decreasing, giving the desired inequality for $\theta^{\prime}>\theta$.
@@
$v(\theta,x)+v(\theta,y)=\int_{0}^{v}[\partial_{1}v(t,x)+\partial_{1}v(t,y)]dt\leq\int_{0}^{v}\partial_{1}v(t,x+y)dt=v(\theta,x+y).$
Now we are ready to prove Theorem 3:
The proof follows the same structure as the proof for the linear valuations. In the following, however, we will assume the stronger notion of Sybil-proofness, that is, the agent can employ an arbitrary finite number of Sybil bids.
@@
$U_{1}(u,v)\geq v\left(\theta_{1},\frac{n-1}{n}\right)-(n-1)v(\theta_{2},\frac{1}{n}).$
Since $v(\theta_{2},\cdot)$ is superadditive, $(n-1)v(\theta_{2},\frac{1}{n})\leq v(\theta_{2},1-/n)$. By continuity of $v$,
@@
which contradicts 8. Therefore $x_{1}(z,\theta)=0$ for all $z\in[0,\theta)$.
In the previous lemma, we have established that for the case of two bidders, the allocation rule always assigns all the item to the highest value bidder. Next, we show by induction on the number of bidders that in the case of more than two bidders, the item is also allocated to the highest value bidder:
We claim that for any $n\geq 2$ and any $\theta_{1}^{\prime}<\theta_{1}:=\max\{\theta_{2},\dots,\theta_{n}\}$, Bidder $1$ will not obtain the object if reports $\theta_{1}^{\prime}$, i.e. $x_{1}(\theta_{1}^{\prime},\theta_{2},\dots,\theta_{n})=0$. We proceed by induction on $n$. We already know the base case $n=2$ is true. Suppose that the claim is true for $n=k$ and assume towards contradiction that $x_{1}(\theta_{1}^{\prime},\theta_{2},\dots,\theta_{k+1})>0$ for some $\theta_{1}^{\prime}<\theta_{1}:=\max\{\theta_{2},\dots.\theta_{k+1}\}$. Since $\theta\mapsto x_{1}(\theta,\theta_{2},\dots,\theta_{k+1})$ is non-decreasing, by Lemma 8 this implies that $\partial_{1}v(z,x_{1}(z,\theta_{2},\dots,\theta_{k+1}))>0$ for all $\theta\geq\theta_{1}^{\prime}$. Consider the scenario where there are $k$ bidders, where Bidder $1$ has value $\theta_{1}$ and Bidder $i$ has value $\theta_{i}$ for $2\leq i\leq k$. If Bidder $1$ bids truthfully, his utility payoff is
@@
by induction hypothesis. However, if Bidder $1$ deviates by bidding $\theta_{1}$ himself and creating a Sybil that bids $\theta_{k+1}$, then his utility payoff is (where $x_{i}(z):=x_{i}(z,\theta_{-i})$ and $p_{i}(z):=p_{i}(z,\theta_{-i})$ for $i=1,k+1$ and $z\in\mathbb{R}_{+}$):

### v5

@@ Appendix C Proof of Theorem 4
If $v$ is strictly increasing in $\mathbb{R}_{>0}\times(0,1]$ and twice differentiable function with $v(\theta,0)=0$ and $v(0,x)=0$ and we have type-convexity and superadditive interim utility then:
@@
(2) Fix $x^{\prime}>x$ and define $\phi(\theta):=v(\theta,x^{\prime})-v(\theta,x)$. Then $\phi^{\prime}(\theta)=\partial_{1}v(\theta,x^{\prime})-\partial_{1}v(\theta,x)$. Because $\partial_{1}v(\theta,0)=0$ and $x\mapsto\partial_{1}v(\theta,x)$ is superadditive, it is non-decreasing on $(0,1]$; hence $\phi^{\prime}(\theta)\geq 0$ for all $\theta>0$. Thus $\phi$ is non-decreasing, giving the desired inequality for $\theta^{\prime}>\theta$.
@@
$v(\theta,x)+v(\theta,y)=\int_{0}^{\theta}[\partial_{1}v(t,x)+\partial_{1}v(t,y)]dt\leq\int_{0}^{\theta}\partial_{1}v(t,x+y)dt=v(\theta,x+y).$
Now we are ready to prove Theorem 4. The proof follows the same structure as the proof for the linear valuations. In the following, however, we will assume the stronger notion of Sybil-proofness, that is, the agent can employ an arbitrary finite number of Sybil bids.
@@
$U_{1}(u,v)\geq v\left(\theta_{1},\frac{n-1}{n}\right)-(n-1)v\left(\theta_{2},\frac{1}{n}\right).$
Since $v(\theta_{2},\cdot)$ is superadditive, $(n-1)v(\theta_{2},\frac{1}{n})\leq v(\theta_{2},1-1/n)$. By continuity of $v$,
@@
which contradicts 8. Therefore, $x_{1}(z,\theta)=0$ for all $z\in[0,\theta)$.
In the previous lemma, we have established that for the case of two bidders, the allocation rule always assigns the whole item to the highest value bidder. Next, we show by induction on the number of bidders that in the case of more than two bidders, the item is also allocated to the highest value bidder.
We claim that for any $n\geq 2$ and any $\theta_{1}^{\prime}<\theta_{1}:=\max\{\theta_{2},\dots,\theta_{n}\}$, Bidder $1$ will not obtain the object if he reports $\theta_{1}^{\prime}$, i.e. $x_{1}(\theta_{1}^{\prime},\theta_{2},\dots,\theta_{n})=0$. We proceed by induction on $n$. We already know the base case $n=2$ is true. Suppose that the claim is true for $n=k$ and assume towards contradiction that $x_{1}(\theta_{1}^{\prime},\theta_{2},\dots,\theta_{k+1})>0$ for some $\theta_{1}^{\prime}<\theta_{1}:=\max\{\theta_{2},\dots.\theta_{k+1}\}$. Assume that the types are labeled so that $\theta_{2}\geq\theta_{3}\geq\dots\geq\theta_{k}\geq\theta_{k+1}$. Since $\theta\mapsto x_{1}(\theta,\theta_{2},\dots,\theta_{k+1})$ is non-decreasing, by Lemma 7 this implies that $\partial_{1}v(z,x_{1}(z,\theta_{2},\dots,\theta_{k+1}))>0$ for all $\theta\geq\theta_{1}^{\prime}$. Consider the scenario where there are $k$ bidders, Bidder $1$ has value $\theta_{1}$ and Bidder $i$ has value $\theta_{i}$ for $2\leq i\leq k$. If Bidder $1$ bids truthfully, his utility payoff is
@@
by induction hypothesis, since $\theta_{1}:=\max(\theta_{2},\dots\theta_{k+1})=\max(\theta_{2},\dots,\theta_{k})$, by labeling. However, if Bidder $1$ deviates by bidding $\theta_{1}$ himself and creating a Sybil that bids $\theta_{k+1}$, then his utility payoff is (where $x_{i}(z):=x_{i}(z,\theta_{-i})$ and $p_{i}(z):=p_{i}(z,\theta_{-i})$ for $i=1,k+1$ and $z\in\mathbb{R}_{+}$):